| : \arrowvert ∥ : \Arrowvert ∥ : \Vert ∥ : \| | : \vert ↑ : \uparrow ↓ : \downarrow ↕ : \updownarrow ⇑ : \Uparrow ⇓ : \Downarrow ⇕ : \Updownarrow \ : \backslash ⟩ : \rangle ⟨ : \langle } : \rbrace { : \lbrace ⌉ : \rceil ⌈ : \lceil ⌋ : \rfloor ⌊ : \lfloor : \lgroup

: \rgroup

: \bracevert ¶ : \mathparagraph § : \mathsection $ : \mathdollar £ : \mathsterling : \mathunderscore …: \mathellipsis ↓ : \downarrow ⇐ : \Leftarrow ⇔ : \Leftrightarrow ⇚ : \Lleftarrow ⇒ : \Rightarrow ⇛ : \Rrightarrow ⟲ : \circlearrowleft ⟳ : \circlearrowright ↶ : \curvearrowleft ↷ : \curvearrowright ⇊ : \downdownarrows ← : \leftarrow ↞ : \twoheadleftarrow ↠ : \twoheadrightarrow ↢ : \leftarrowtail ↣ : \rightarrowtail ↪ : \hookrightarrow ↩ : \hookleftarrow ⇇ : \leftleftarrows ↔ : \leftrightarrow ⇆ : \leftrightarrows ↫ : \looparrowleft ↬ : \looparrowright ⇍ : \nLeftarrow ⇎ : \nLeftrightarrow ⇏ : \nRightarrow ↗ : \nearrow ↚ : \nleftarrow ↮ : \nleftrightarrow ↛ : \nrightarrow ↖ : \nwarrow → : \rightarrow ⇄ : \rightleftarrows ⇉ : \rightrightarrows ↭ : \leftrightsquigarrow ⇝ : \rightsquigarrow ↘ : \searrow ↙ : \swarrow ⇈ : \upuparrows ▷◁ : \bowtie ⊧ : \models ⟹ : \Longrightarrow ⟶ : \longrightarrow ⟵ : \longleftarrow ⟸ : \Longleftarrow ⟼ : \longmapsto ⟷ : \longleftrightarrow ⟺ : \Longleftrightarrow ⟺ : \iff ⋯: \cdots

⋮ : \vdots

... : \ddots = : \Relbar − : \relbar ≐ : \doteq ⇌ : \rightleftharpoons ∉ : \notin ≅ : \cong ↦ : \mapsto → : \to ← : \gets ∋ : \owns ≥ : \ge ≤ : \le ≠ : \neq ∨ : \lor ∧ : \land ∮ : \oint ∫ : \int ¬ : \lnot ∠ : \angle √ : \surd ℏ : \hbar \overrightarrow{x} : \overrightarrow x \overleftarrow{x} : \overleftarrow x \overbrace{x} : \overbrace x \underbrace{x} : \underbrace x bc : \skew 3 b c ⟶: \rightarrowfill ⟵: \leftarrowfill : \downbracefill : \upbracefill ∥ : \\ | £ : \mathsterling : \mathunderscore …: \mathellipsis ℓ : \ell Γ : \Gamma Δ : \Delta Λ : \Lambda Ω : \Omega Φ : \Phi Π : \Pi Ψ : \Psi Σ : \Sigma Θ : \Theta Υ : \Upsilon Ξ : \Xi α : \alpha β : \beta χ : \chi δ : \delta ϵ : \epsilon η : \eta γ : \gamma ι : \iota κ : \kappa λ : \lambda μ : \mu ν : \nu ω : \omega ϕ : \phi π : \pi ψ : \psi ρ : \rho σ : \sigma τ : \tau θ : \theta υ : \upsilon ε : \varepsilon κ : \varkappa φ : \varphi ϖ : \varpi ϱ : \varrho ς : \varsigma ϑ : \vartheta ξ : \xi ζ : \zeta ℵ : \aleph ℶ : \beth ℸ : \daleth ⇃ : \downharpoonleft ⇂ : \downharpoonright ↽ : \leftharpoondown ↼ : \leftharpoonup ⇋ : \leftrightharpoons ⇁ : \rightharpoondown ⇀ : \rightharpoonup ↿ : \upharpoonleft ↾ : \upharpoonright ♣ : \clubsuit ♢ : \diamondsuit ♡ : \heartsuit ♠ : \spadesuit 𝕜 : \Bbbk ≎ : \Bumpeq ⋒ : \Cap ⋓ : \Cup Ⅎ : \Finv ⅁ : \Game ℑ : \Im ↰ : \Lsh ℜ : \Re ↱ : \Rsh ⋐ : \Subset ⋑ : \Supset ⊩ : \Vdash ⊪ : \Vvdash ⨿ : \amalg ≈ : \approx ≊ : \approxeq

∗ : \ast ≍ : \asymp ϶ : \backepsilon ‵ : \backprime ∽ : \backsim ⋍ : \backsimeq ⊼ : \barwedge ∵ : \because ≬ : \between ⋂ : \bigcap ◯ : \bigcirc ⋃ : \bigcup ⨀ : \bigodot ⨁ : \bigoplus ⨂ : \bigotimes ⨆ : \bigsqcup ⋆ : \bigstar ▽ : \bigtriangledown △ : \bigtriangleup ⨄ : \biguplus ⋁ : \bigvee ⋀ : \bigwedge ♦ : \blacklozenge ■ : \blacksquare ▲ : \blacktriangle ▼ : \blacktriangledown ◀ : \blacktriangleleft ▶ : \blacktriangleright ⊥ : \bot ⊡ : \boxdot ⊟ : \boxminus ⊞ : \boxplus ⊠ : \boxtimes : \braceld : \bracelu : \bracerd : \braceru

• : \bullet ≏ : \bumpeq ∩ : \cap · : \cdot · : \cdotp ▪ : \centerdot ∘ : \circ ≗ : \circeq Ⓢ : \circledS ⊛ : \circledast ⊚ : \circledcirc ⊝ : \circleddash : : \colon ∁ : \complement ∐ : \coprod ∪ : \cup ⋞ : \curlyeqprec ⋟ : \curlyeqsucc ⋎ : \curlyvee ⋏ : \curlywedge † : \dagger ⊣ : \dashv ‡ : \ddagger ⧹ : \diagdown ⧸ : \diagup ⋄ : \diamond 𝟋 : \digamma ÷ : \div ⋇ : \divideontimes ≑ : \doteqdot ∔ : \dotplus ⩞ : \doublebarwedge ∅ : \emptyset ≖ : \eqcirc ≂ : \eqsim ⪖ : \eqslantgtr ⪕ : \eqslantless ≡ : \equiv ð : \eth ∃ : \exists ≒ : \fallingdotseq ♭ : \flat ∀ : \forall ⌢ : \frown ≥ : \geq ≧ : \geqq ⩾ : \geqslant ≫ : \gg ≫ : \ggg ℷ : \gimel ⪊ : \gnapprox ⪈ : \gneq ≩ : \gneqq ⋧ : \gnsim ⪆ : \gtrapprox ⋗ : \gtrdot ⋛ : \gtreqless ⪌ : \gtreqqless ≷ : \gtrless ≳ : \gtrsim ≩ : \gvertneqq ℏ : \hslash ı : \imath ∈ : \in ∞ : \infty ⊺ : \intercal ∫ : \intop ȷ : \jmath . : \ldotp ⋋ : \leftthreetimes ≤ : \leq ≦ : \leqq ⩽ : \leqslant ⪅ : \lessapprox ⋖ : \lessdot ⋚ : \lesseqgtr ⪋ : \lesseqqgtr ≶ : \lessgtr ≲ : \lesssim : \lhook ≪ : \ll ≪ : \lll ⪉ : \lnapprox ⪇ : \lneq ≨ : \lneqq ⋦ : \lnsim ♢ : \lozenge ⋉ : \ltimes ≨ : \lvertneqq : \mapstochar $ : \mathdollar ¶ : \mathparagraph § : \mathsection ∡ : \measuredangle | : \mid ∓ : \mp ⊸ : \multimap ⊯ : \nVDash ⊮ : \nVdash ∇ : \nabla ♮ : \natural ≇ : \ncong ¬ : \neg ∄ : \nexists ≱ : \ngeq ≧̸ : \ngeqq ⩾̸ : \ngeqslant ≯ : \ngtr ∋ : \ni ≰ : \nleq ≦̸ : \nleqq ⩽̸ : \nleqslant ≮ : \nless ∤ : \nmid : \not ∦ : \nparallel ⊀ : \nprec ⪯̸ : \npreceq ∤ : \nshortmid ∦ : \nshortparallel ≁ : \nsim ⊈ : \nsubseteq ⫅̸ : \nsubseteqq ⊁ : \nsucc ⪰̸ : \nsucceq ⊉ : \nsupseteq ⫆̸ : \nsupseteqq ⋪ : \ntriangleleft ⋬ : \ntrianglelefteq ⋫ : \ntriangleright ⋭ : \ntrianglerighteq ⊭ : \nvDash ⊬ : \nvdash ⊙ : \odot ∮ : \ointop ⊖ : \ominus ⊕ : \oplus ⊘ : \oslash ⊗ : \otimes ∥ : \parallel ∂ : \partial ⊥ : \perp ⋔ : \pitchfork ± : \pm ≺ : \prec ≾ : \precapprox ≼ : \preccurlyeq ⪯ : \preceq ⪹ : \precnapprox ⪵ : \precneqq ⋨ : \precnsim ≾ : \precsim ′ : \prime ∏ : \prod ∝ : \propto : \rhook ⋌ : \rightthreetimes ≓ : \risingdotseq ⋊ : \rtimes \ : \setminus ♯ : \sharp ∣ : \shortmid ∥ : \shortparallel ∼ : \sim ≃ : \simeq ⌢ : \smallfrown ∫ : \smallint ∖ : \smallsetminus ⌣ : \smallsmile ⌣ : \smile ∢ : \sphericalangle ⊓ : \sqcap ⊔ : \sqcup ⊑ : \sqsubseteq ⊒ : \sqsupseteq □ : \square ⋆ : \star ⊂ : \subset ⊆ : \subseteq ⫅ : \subseteqq ⊊ : \subsetneq ⫋ : \subsetneqq ≻ : \succ ≿ : \succapprox ≽ : \succcurlyeq ⪰ : \succeq ⪺ : \succnapprox ⪶ : \succneqq ⋩ : \succnsim ≿ : \succsim ∑ : \sum ⊃ : \supset ⊇ : \supseteq ⫆ : \supseteqq ⊋ : \supsetneq ⫌ : \supsetneqq ∴ : \therefore ≈ : \thickapprox ∼ : \thicksim × : \times ⊤ : \top △ : \triangle ▽ : \triangledown ◁ : \triangleleft ⊴ : \trianglelefteq ≜ : \triangleq ▷ : \triangleright ⊵ : \trianglerighteq ⊎ : \uplus ⊨ : \vDash ∅ : \varnothing ∝ : \varpropto ⊊ : \varsubsetneq ⫋ : \varsubsetneqq ⊋ : \varsupsetneq ⫌ : \varsupsetneqq △ : \vartriangle ◁ : \vartriangleleft ▷ : \vartriangleright ⊢ : \vdash ∨ : \vee ⊻ : \veebar ∧ : \wedge ℘ : \wp ≀ : \wr